OpenStudy (anonymous):

The expression below shows the population of foxes in a forest after 12x years: f(x) = 50(1.05)^12x Which of the following is an equivalent function? f(x) = 50(1.7959)^x f(x) = 50(12.6000)^x f(x) = 600(1.05)^x f(x) = 4.2(1.05)^x

OpenStudy (anonymous):

@$MilitaryLad$

OpenStudy (anonymous):

It should be equivalent to 50(1.7959)^x, rather very very similar

OpenStudy (bibby):

\({a^bc}=(a^b)^c\) so here (1.05)^12x=(1.05^12)x

OpenStudy (anonymous):

Its C

OpenStudy (bibby):

(1.05^12)^x*

OpenStudy (anonymous):

who was right?

OpenStudy (bibby):

you

OpenStudy (anonymous):

So, A, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

tank yew!

OpenStudy (anonymous):

C is the best guess

OpenStudy (bibby):

it's A. like Apple

OpenStudy (anonymous):

who was right?